# Evaluating Optimal Reference Translations

### 28 Nov 2023 v1

@@ 1 Introduction
After discussing related work in this context (Section 2), we focus on defining ORT and describe its creation process (Section 3). Next, we describe our evaluation campaign of ORT, the data, annotation interface, and annotation instructions (Section 3.2). We then turn to a statistical perspective of our data and measure the predictability of human ratings (e.g. Overall rating from Spelling, Style, Meaning, etc.) using automated metrics (Section 4). We pay special attention to predicting document-level rating from segment-level. In the penultimate Section 5, we provide a detailed qualitative analysis of human annotations and discuss this work in the greater perspective of human evaluation of translations (Section 6). Analysis code and collected data are publicly available.222github.com/ufal/optimal-reference-translations
@@ 3.1 Translation Creation
The underlying dataset without the evaluation has already been described in Czech (Kloudová et al., , 2022). The 130 original English texts (news articles available from the Internet, covering topics ranging from politics and economics to sports and social events) were translated from English into Czech by three human translators for the Conference on Machine Translation 2020 (WMT20). The three translators were hired by WMT organizers from a translation agency. The resulting three independent parallel Czech translations (P1, P2, P3) serve as basic reference translations, from which a final “optimal reference translation” could be synthesized. It was anticipated that our creators of ORT (two translators-cum-theoreticians – professionals who deal with translation from both a practical and a theoretical point of view)333One of them is a co-author of this article. However, the translations are later independently evaluated and hence, to the best of our knowledge and conscience, we do not consider this to be a conflict of interest or otherwise a methodological flaw. would always choose the best translation solutions from the existing three versions, or create new solutions if necessary. However, the available translations from the first-stage translators were often of insufficient quality. Therefore, in the creation of our optimal reference translations, more emphasis was placed on the input of the creators of the final version rather than on the synthesis of existing translations.
@@ 3.2.1 Annotators
We hired 11 native Czech annotators for the evaluation of translations in three groups: (1) four professional translators,444Defining who is a professional translator is not easy. The factors influencing the degree of professionalism of a translator include, among others, education and experience. Professional translators in our study have at least one of the following: (1) completed an M.A. degree programme in English-Czech translation studies, (2) completed an M.A. degree programme in interpreting or philology, or (3) have at least 10 years of translation experience. (2) four non-experts, (3) three students of MA Study Programme Translation and/or Interpreting: Czech and English at the Institute for Translation Studies.555utrl.ff.cuni.cz Their proficiency and end-campaign questionnaire responses are presented in Section 4.1.
@@ 4.1 Annotator Questionnaire
We did not constrain the annotators in what order they should perform the annotations. As a result, they employed various approaches, most popular being segment-category-translation.666I.e. first finish all annotation categories in a translation, then all annotation categories in the second translation, etc., and afterwards move to the second segment. While we attempted to not introduce a bias, almost all annotators filled in categories one-by-one as they were organized in the user interface.777I.e. starting from Spelling and ending with Overall. This could have an effect on the rating. For example, by establishing and drawing attention to the specific 6 features, the final Overall rating may be influenced primarily by them and it would not have been if the ordering was reversed. Pragmatics and Overall were reported as the hardest to evaluate, while Spelling was the easiest, especially because errors in spelling can be seen even without deeper translatological analysis and there were not many of them in the translations. The annotators self-reported utilizing the preceding and following context around half the time to check for document-level consistency. While they proceeded mostly linearly, about 20% (self-reported estimate) of previously completed segments were later changed. We intentionally shuffled the ordering of translations (columns in each sheet) so that the annotators would not build a bias towards the translation source in e.g. the second column. However, the annotators reported that despite this, they were sometimes able to recognize a specific translation source based on various artifacts, such as systematically not translating or localizing foreign names.
@@ 4.4 Inter-Annotator Agreement
To measure inter-annotator agreement, we aggregate pairwise annotator Pearson correlations on the segment-level.888Even though the data are not normally distributed, the Perason correlation reveals agreement controlled for each annotator’s mean and variance. At first, this agreement is quite low ( = 0.33). It can however be explained upon closer inspection of agreement across translations. While inter-annotator correlations for the the worst translation P3 were = 0.50, the best translation had = 0.13. We hypothesize that with less variance and therefore signal for rating, the inter-annotator agreement drops. This is even more visible from the pairwise annotator correlations for the Grammar category, in which N1 has made almost no errors ( = 0.03). In 28% of cases, the ordering of Overall scores for segments was the same between pairs of annotators and in 66% of cases they differed by only one transposition. In other words, the difference in the score ordering was 2 positions or more only in 8% of cases. Further individual effects of annotators are discussed in Section 4.6.
@@ 4.5.2 Linear Regression on Other Categories
We treat the prediction of Overall from other categories as a regression task with 6 numerical input features (Spelling, Terminology, etc) and one numerical output feature (Overall). We subtract the mean to preserve only the variance to be able to interpret the learned coefficients of a linear regression model. We split document- and segment-level ratings into train/test as 778/100 and 6925/100, respectively. Figure 6 shows the results of fitting two linear regression models together with the coefficients of individual variables. Because the distributions of features is similar, as documented in Figure 3, we can interpret the magnitude of the coefficient as the importance in determining the Overall score. For both the document and segment level, Spelling and Meaning have the highest impact while Terminology and Style have the least impact.999These interpretations are, however, not fully conclusive because of a possible latent co-dependent. The Overall variable may in reality be largely dependent on another variable for which we do not have annotations. One hypothetical translation source could be very good if measured on the variable and also Overall and unrelated to that also good in the Spelling level, which would yield similar results to those presented. The linear regression model is further negatively affected by the non-linearity of the human bias towards round numbers, which the model is not able to take into consideration. The fitted coefficients are confirmed by annotator reponses in the questionnaire in which Meaning, Style and Pragmatics was most important to them when evaluating Overall.
@@ 4.5.3 Automated Metrics
As mentioned in Section 3.2, annotators were tasked to post-edit texts to a state which they would be content with. As a result, the annotators post-edited 62% of all the segments on average. We compute several automatic metric scores between the original and edited versions of segments and compare them to the collected scores, such as Overall. This allows us to answer the question: Does the post-edited distance (as measured by automated metrics) correspond to the annotator score (negatively)? The results in Footnote 10 show that there is very little difference between individual metrics. Most score categories are equally predictive with the exception of Overall (most) and Spelling (least). The explanation for this phenomena for Spelling is again (Section 4.4) much lower variance. Overall, the more the annotators changed the original text in their post-editing, the lower score they assigned to the hypothesis. Including the metrics in the prediction of Overall in Section 4.5.2 does not provide any additional improvement on top of other categories (final segment-level is still ).
@@ 4.6 Annotator Differences
Recall that we considered three types of annotators: professional translators, students of translation and non-translators. Despite the same annotation guidelines, their approach to the task was vastly different. For example, Figure 8 shows the distribution of segment-level ratings of Overall. Professional translators produced much more varying and spread-out distribution, especially compared to non-translators, who rated most segments very high. The group differences should be taken into account when modelling the annotation process statistically. When predicting segment-level Overall from other categories, as in Section 4.5.2, the individual annotator Pearson correlation ranges from as high as to as low as . Similar to results of Karpinska et al., (2021) we find that expert annotators are important and have less noise. The average correlations with Overall for the translator, student and non-translator groups are , and , respectively. The expertise feature alone yields correlation with Overall and users alone . This is expected as the groups and users have different means of the variable. This information can be used in combination with other predictive features to push the segment-level correlation from (Figure 6) to . Greater improvement is achieved when combined with the editing distance, such as pushing BLEU from (Footnote 10) to when individual annotators are considered as an input feature (one-hot encoded).
@@ 5.3 Document-level Phenomena
Documents in which certain terms occur that should be consistent throughout the text and/or should correspond in a meaningful way to the thematic and pragmatic area, appear to be appropriate material to demonstrate this (spans marked with 1).111111We mark all related spans in the source as well as in all discussed translations even if they are correct, for easy comparison. These observations are in line with MT evaluation methods focused on terminology (Zouhar et al., , 2020; Semenov and Bojar, , 2022; Agarwal et al., , 2023). Another phenomenon to be observed might be a particular way of spelling words, which generally have two or more accepted spellings and the convention is just to achieve a consistent spelling throughout the document (spans marked with 2).
Finally, for the topic-focus articulation, also called functional sentence perspective (spans marked with 3), it is crucial to respect the context of the whole document (Daneš, , 1974; Sgall et al., , 1986; Hajičová et al., , 2013). It is concerned with the distribution of information as determined by all meaningful elements, including context. In Section 5.3, evaluated by a non-translator, we selectively document the distribution of the degrees of communicative dynamism over sentence elements in Czech, which determines the orientation or perspective of the sentence (Firbas, , 1992).121212The examples in Section 5.3 also contain other translation errors, such as incorrect name translation Grubb/Brubb and not only those related to the document-level phenomena. Unless otherwise stated, the evaluator has left these errors in the translation even after editing. Since we present associated discussions in more detail within the previous sections, we do not elaborate on them at this point.
@@ 7 Conclusion
While we focused on evaluating human translations, the idential setup could be used for evaluating MT models, which we plan to address in future work. This is not part of the present work which is focused on showing that the reference translations usually used are of insufficient quality and need to be reconsidered. Our next step will be to assess which of the multitude of automatic metrics of MT quality are sensitive to the subtleties captured in our ORT and can thus be used to reliably evaluate MT outputs of high quality. This will again require careful expert manual evaluation.

### 08 Mar 2024 v2

@@ 1 Introduction
After discussing related work in this context (Section 2), we focus on defining ORT and describe its creation process (Section 3). Next, we describe our evaluation campaign of ORT, the data, annotation interface, and annotation instructions (Section 3.2). We then turn to a statistical perspective of our data and measure the predictability of human ratings (e.g. Overall rating from Spelling, Style, Meaning, etc.) using automated metrics (Section 4). We pay special attention to predicting document-level rating from segment-level. In the penultimate Section 5, we provide a detailed qualitative analysis of human annotations and discuss this work in the greater perspective of human evaluation of translations (Section 6). Analysis code and collected data are publicly available.
@@ 3.1 Translation Creation
The underlying dataset without the evaluation has already been described in Czech (Kloudová et al., , 2022). The 130 original English texts (news articles available from the Internet, covering topics ranging from politics and economics to sports and social events) were translated from English into Czech by three human translators for the Conference on Machine Translation 2020 (WMT20). The three translators were hired by WMT organizers from a translation agency. The resulting three independent parallel Czech translations (P1, P2, P3) serve as basic reference translations, from which a final “optimal reference translation” could be synthesized. It was anticipated that our creators of ORT (two translators-cum-theoreticians – professionals who deal with translation from both a practical and a theoretical point of view)222One of them is a co-author of this article. However, the translations are later independently evaluated and hence, to the best of our knowledge and conscience, we do not consider this to be a conflict of interest or otherwise a methodological flaw. would always choose the best translation solutions from the existing three versions, or create new solutions if necessary. However, the available translations from the first-stage translators were often of insufficient quality. Therefore, in the creation of our optimal reference translations, more emphasis was placed on the input of the creators of the final version rather than on the synthesis of existing translations.
@@ 3.2.1 Annotators
We hired 11 native Czech annotators for the evaluation of translations in three groups: (1) four professional translators,333Defining who is a professional translator is not easy. The factors influencing the degree of professionalism of a translator include, among others, education and experience. Professional translators in our study have at least one of the following: (1) completed an M.A. degree programme in English-Czech translation studies, (2) completed an M.A. degree programme in interpreting or philology, or (3) have at least 10 years of translation experience. (2) four non-experts, (3) three students of MA Study Programme Translation and/or Interpreting: Czech and English at the Institute for Translation Studies.444utrl.ff.cuni.cz Their proficiency and end-campaign questionnaire responses are presented in Section 4.1.
@@ 4.1 Annotator Questionnaire
We did not constrain the annotators in what order they should perform the annotations. As a result, they employed various approaches, most popular being segment-category-translation.555I.e. first finish all annotation categories in a translation, then all annotation categories in the second translation, etc., and afterwards move to the second segment. While we attempted to not introduce a bias, almost all annotators filled in categories one-by-one as they were organized in the user interface.666I.e. starting from Spelling and ending with Overall. This could have an effect on the rating. For example, by establishing and drawing attention to the specific 6 features, the final Overall rating may be influenced primarily by them and it would not have been if the ordering was reversed. Pragmatics and Overall were reported as the hardest to evaluate, while Spelling was the easiest, especially because errors in spelling can be seen even without deeper translatological analysis and there were not many of them in the translations. The annotators self-reported utilizing the preceding and following context around half the time to check for document-level consistency. While they proceeded mostly linearly, about 20% (self-reported estimate) of previously completed segments were later changed. We intentionally shuffled the ordering of translations (columns in each sheet) so that the annotators would not build a bias towards the translation source in e.g. the second column. However, the annotators reported that despite this, they were sometimes able to recognize a specific translation source based on various artifacts, such as systematically not translating or localizing foreign names.
@@ 4.4 Inter-Annotator Agreement
To measure inter-annotator agreement, we aggregate pairwise annotator Pearson correlations on the segment-level.777Even though the data are not normally distributed, the Perason correlation reveals agreement controlled for each annotator’s mean and variance. At first, this agreement is quite low ( = 0.33). It can however be explained upon closer inspection of agreement across translations. While inter-annotator correlations for the the worst translation P3 were = 0.50, the best translation had = 0.13. We hypothesize that with less variance and therefore signal for rating, the inter-annotator agreement drops. This is even more visible from the pairwise annotator correlations for the Grammar category, in which N1 has made almost no errors ( = 0.03). In 28% of cases, the ordering of Overall scores for segments was the same between pairs of annotators and in 66% of cases they differed by only one transposition. In other words, the difference in the score ordering was 2 positions or more only in 8% of cases. Further individual effects of annotators are discussed in Section 4.6.
@@ 4.5.2 Linear Regression on Other Categories
We treat the prediction of Overall from other categories as a regression task with 6 numerical input features (Spelling, Terminology, etc) and one numerical output feature (Overall). We subtract the mean to preserve only the variance to be able to interpret the learned coefficients of a linear regression model. We split document- and segment-level ratings into train/test as 778/100 and 6925/100, respectively. Figure 6 shows the results of fitting two linear regression models together with the coefficients of individual variables. Because the distributions of features is similar, as documented in Figure 3, we can interpret the magnitude of the coefficient as the importance in determining the Overall score. For both the document and segment level, Spelling and Meaning have the highest impact while Terminology and Style have the least impact.888These interpretations are, however, not fully conclusive because of a possible latent co-dependent. The Overall variable may in reality be largely dependent on another variable for which we do not have annotations. One hypothetical translation source could be very good if measured on the variable and also Overall and unrelated to that also good in the Spelling level, which would yield similar results to those presented. The linear regression model is further negatively affected by the non-linearity of the human bias towards round numbers, which the model is not able to take into consideration. The fitted coefficients are confirmed by annotator reponses in the questionnaire in which Meaning, Style and Pragmatics was most important to them when evaluating Overall.
@@ 4.5.3 Automated Metrics
As mentioned in Section 3.2, annotators were tasked to post-edit texts to a state which they would be content with. As a result, the annotators post-edited 62% of all the segments on average. We compute several automatic metric scores between the original and edited versions of segments and compare them to the collected scores, such as Overall. This allows us to answer the question: Does the post-edited distance (as measured by automated metrics) correspond to the annotator score (negatively)? The results in Footnote 9 show that there is very little difference between individual metrics. Most score categories are equally predictive with the exception of Overall (most) and Spelling (least). The explanation for this phenomena for Spelling is again (Section 4.4) much lower variance. Overall, the more the annotators changed the original text in their post-editing, the lower score they assigned to the hypothesis. Including the metrics in the prediction of Overall in Section 4.5.2 does not provide any additional improvement on top of other categories (final segment-level is still ).
@@ 4.6 Annotator Differences
Recall that we considered three types of annotators: professional translators, students of translation and non-translators. Despite the same annotation guidelines, their approach to the task was vastly different. For example, Figure 8 shows the distribution of segment-level ratings of Overall. Professional translators produced much more varying and spread-out distribution, especially compared to non-translators, who rated most segments very high. The group differences should be taken into account when modelling the annotation process statistically. When predicting segment-level Overall from other categories, as in Section 4.5.2, the individual annotator Pearson correlation ranges from as high as to as low as . Similar to results of Karpinska et al., (2021) we find that expert annotators are important and have less noise. The average correlations with Overall for the translator, student and non-translator groups are , and , respectively. The expertise feature alone yields correlation with Overall and users alone . This is expected as the groups and users have different means of the variable. This information can be used in combination with other predictive features to push the segment-level correlation from (Figure 6) to . Greater improvement is achieved when combined with the editing distance, such as pushing BLEU from (Footnote 9) to when individual annotators are considered as an input feature (one-hot encoded).
@@ 5.3 Document-level Phenomena
Documents in which certain terms occur that should be consistent throughout the text and/or should correspond in a meaningful way to the thematic and pragmatic area, appear to be appropriate material to demonstrate this (spans marked with 1).101010We mark all related spans in the source as well as in all discussed translations even if they are correct, for easy comparison. These observations are in line with MT evaluation methods focused on terminology (Zouhar et al., , 2020; Semenov and Bojar, , 2022; Agarwal et al., , 2023). Another phenomenon to be observed might be a particular way of spelling words, which generally have two or more accepted spellings and the convention is just to achieve a consistent spelling throughout the document (spans marked with 2).
Finally, for the topic-focus articulation, also called functional sentence perspective (spans marked with 3), it is crucial to respect the context of the whole document (Daneš, , 1974; Sgall et al., , 1986; Hajičová et al., , 2013). It is concerned with the distribution of information as determined by all meaningful elements, including context. In Section 5.3, evaluated by a non-translator, we selectively document the distribution of the degrees of communicative dynamism over sentence elements in Czech, which determines the orientation or perspective of the sentence (Firbas, , 1992).111111The examples in Section 5.3 also contain other translation errors, such as incorrect name translation Grubb/Brubb and not only those related to the document-level phenomena. Unless otherwise stated, the evaluator has left these errors in the translation even after editing. Since we present associated discussions in more detail within the previous sections, we do not elaborate on them at this point.
@@ 7 Conclusion
While we focused on evaluating human translations, the idential setup could be used for evaluating MT models, which we plan to address in future work (Zouhar and Bojar, , 2024). This is not part of the present work which is focused on showing that the reference translations usually used are of insufficient quality and need to be reconsidered. Our next step will be to assess which of the multitude of automatic metrics of MT quality are sensitive to the subtleties captured in our ORT and can thus be used to reliably evaluate MT outputs of high quality. This will again require careful expert manual evaluation.